a) Use the constant voltage drop model for the diode and determine the input / output ratio (Vout x Vin), equations. Vdon = 0.7V, rz = 0Ω, Vz = 3V and R = R1 = R2 = 1000Ω.

b) Draw the Vout x Vin graph. Highlight the slope coefficients of the lines and the coordinates of the points where the slope changes occur.

D1 D2 DR3 Vin 7R1 Vout 214 22
Answer #1

circuit as. (-ve input) for analysis change the below. i ☆ t Vin 1 R2 Vout R for will be open (Vout= vin -3 <vin <o diode curD2 will be on. for voltages less than -4.4 Vout -3-007 You (out = -3-7 for ove axis graph a Yout -4.4 -BU - Vn Vin -3.7 2 slo